# TorchCP: A Python Library for Conformal Prediction

### v2

@@ Abstract
Conformal Prediction (CP) has attracted great attention from the research community due to its strict theoretical guarantees. However, researchers and developers still face challenges of applicability and efficiency when applying CP algorithms to deep learning models. In this paper, we introduce TorchCP, a comprehensive PyTorch-based toolkit to strengthen the usability of CP for deep learning models. TorchCP implements a wide range of post-hoc and training methods of conformal prediction for various machine learning tasks, including classification, regression, GNN, and LLM. Moreover, we provide user-friendly interfaces and extensive evaluations to easily integrate CP algorithms into specific tasks. Our TorchCP toolkit, built entirely with PyTorch, enables high-performance GPU acceleration for deep learning models and mini-batch computation on large-scale datasets. With the LGPL license, the code is open-sourced at https://github.com/ml-stat-Sustech/TorchCP and will be continuously updated.
Keywords: Conformal Prediction, Uncertainty Quantification, Machine Learning
@@ 1 Introduction
Conformal prediction (CP) is a powerful statistical framework that provides reliable measures of uncertainty for machine learning predictions. With conformal prediction, the model yields prediction sets or intervals designed to contain the ground truths with a specified coverage probability  (Vovk et al., 2005; Shafer and Vovk, 2008; Balasubramanian et al., 2014; Angelopoulos et al., 2020; Manokhin, 2022). In recent years, conformal prediction algorithms have increasingly been integrated into deep learning models, enabling trustworthy predictions for high-stake scenarios in the real world. The trend raises new challenges of applicability and efficiency in implementing conformal prediction algorithms across various tasks of deep learning, such as time series, graph neural networks, and large language models (Zargarbashi et al., 2023; De Toni et al., 2024; Quach et al., 2024; Dey et al., 2024).
Given the increasing popularity of deep learning, it is indispensable to develop an extensive toolkit to enable easy, fast, and flexible integration of CP algorithms in various deep learning applications. There are existing libraries of uncertainty quantification that contain implementations of some CP algorithms, such as Puncc (Mendil et al., 2023), crepes(Boström, 2022), and Fortuna (Detommaso et al., 2024). However, those implementations are typically limited in the scope of application, only supporting classification and regression tasks. It is essential to develop a comprehensive CP library that supports a wider range of deep learning tasks, such as graph node classification, ordinal classification, and text generation with large language models. Besides, most of the current libraries lack support for GPU acceleration, which is crucial for deep learning models, and do not facilitate mini-batch computation for large datasets. This leads to suboptimal computational efficiency and scalability, particularly for large-scale applications. We summarize a feature comparison of current libraries in Table 1.
Table 1: Open-source libraries with support of conformal prediction.
Library Task Support
Classification Regression Graph LLM Training GPU Mini-Batch Backend
nonconformist [^1] https://github.com/donlnz/nonconformist. Numpy
TorchUQ [^2] https://github.com/TorchUQ/torchuq?tab=readme-ov-file. PyTorch
@@
In this paper, we present TorchCP, a Python library of conformal prediction for deep learning, built on PyTorch with strong GPU acceleration. TorchCP addresses the above limitations through the following crucial features: 1) Comprehensiveness. TorchCP implements a wide range of state-of-the-art CP algorithms, encompassing both post-hoc and training methods. In addition to conventional classification and regression tasks, the framework extends its functionality to various deep learning scenarios, including graph node classification, large language models, and ordinal classification, making it a versatile toolkit for modern machine learning applications. 2) Scalability. TorchCP is designed in a modular fashion and provides user-friendly interfaces for CP algorithms. Researchers can integrate and experiment with various CP algorithms through modular components. The framework’s plug-in architecture also enables the application of CP training algorithms to any existing loss function, offering exceptional flexibility for machine learning practitioners. 3) Efficiency. TorchCP delivers high-performance CP algorithms with strong GPU acceleration. This native PyTorch implementation enables mini-batch computation with high efficiency, providing seamless integration with deep learning models.
In this project, we open-source TorchCP, accompanied by detailed documentation and tutorials, to facilitate the easy, flexible, and efficient implementation of CP algorithms in deep learning applications. We believe that TorchCP will promote the adoption and advancement of conformal prediction within the deep learning community.
Figure 1: The scope of TorchCP.
2 The Architecture Design of TorchCP
2.1 Overview of Implementations
TorchCP implements a wide range of state-of-the-art conformal prediction algorithms, supporting both training approaches and post-hoc methods for various deep learning tasks (Figure 1). Specifically, for classification tasks, TorchCP provides various non-conformity score functions (Gammerman and Vovk, 2007; Löfström et al., 2015; Sadinle et al., 2019; Romano et al., 2020; Angelopoulos et al., 2020; Huang et al., 2024), CP algorithms (Vovk, 2012; Tibshirani et al., 2019; Ding et al., 2024), and conformal trainer  (Stutz et al., 2022; Dey et al., 2024; Xi et al., 2024; Liu et al., 2024). For regression tasks, TorchCP implements the standard conformal regression algorithm (Lei et al., 2018), multiple variants of Conformal Quantile Regression (CQR) (Romano et al., 2019; Kivaranovic et al., 2020; Sesia and Candès, 2020), and various conformal algorithms for time series problems (Gibbs and Candes, 2021; Xu and Xie, 2021).
Moreover, TorchCP provides conformal prediction algorithms for node classification of graphs and large language models (LLMs). Specifically, TorchCP implements various variants of APS that are specifically designed for graph structure (Wijegunawardana et al., 2020; Clarkson, 2023; Zargarbashi et al., 2023; Huang et al., 2023; Song et al., 2024), and the certified risk control algorithm for text generation task of LLMs (Quach et al., 2024).
Finally, to facilitate the evaluation and selection of CP methods, TorchCP offers extensive metrics for assessing the marginal/conditional coverage and size efficiency of prediction sets produced by different CP methods. This enables researchers to benchmark the superiority of novel approaches and identify the most suitable existing methods for specific tasks.
2.2 Modular Design
To improve extensibility and flexibility, TorchCP adopts a highly modularized architecture in implementing conformal prediction methods. For each task, TorchCP implements three main components of conformal prediction algorithms: Trainer, Score and Predictor. We illustrate the modular design of TorchCP on the left of Figure 2. The details of the three components are introduced below.
  • Trainer module defines those training algorithms with loss functions for improving conformal prediction, like ConfTr (Stutz et al., 2022) and C-Adapter (Liu et al., 2024).
  • Score module defines those non-conformity scores that quantify how different a new sample is from the training data, e.g., THR (Sadinle et al., 2019) ,APS (Gibbs and Candes, 2021), SAPS (Huang et al., 2024) and CQR (Romano et al., 2019).
  • Predictor module defines the conformal prediction procedure, including the functions of calibration and evaluation. For example, two common-used conformal predictors are Split CP (Vovk et al., 2005) and Class-wise CP (Vovk, 2012). A conformal predictor is defined with a score and a trained model from the trainer. The evaluate function outputs the two measures of predictors: average set size and empirical coverage rate. We provide more metrics for measuring the predictors in utils/metrics.
Example Code for Conformal Prediction. On the right side of Figure 2, we present an example code snippet for multi-class classification. First, we prepare a calibration dataset and a test dataset (Lines 5-6), along with a pre-trained PyTorch model (Lines 9-10). Next, we define a split predictor using the THR score and the trained model (Line 14). The calibration procedure is conducted by running the calibrate function with the cal_dataloader and the significance level $\alpha$ (Line 17). Finally, we obtain the evaluation results through the evaluate function (Lines 20-21). This modular design is compatible with various deep learning tasks, allowing for easy implementation of new conformal prediction algorithms.
Figure 2: An illustration of the TorchCP framework.
@@ 3 Conclusion and Future work
This paper presents TorchCP, a Python toolbox for conformal prediction research on deep learning models, built on the PyTorch Library with strong GPU acceleration. TorchCP implements a wide range of state-of-the-art conformal prediction algorithms and provides user-friendly interfaces along with comprehensive evaluation functions. Besides, we provide extensive documents, tutorials, and unit tests (100% test coverage) to improve the usability of TorchCP. The toolbox can support future research and applications of conformal prediction algorithms on deep learning models. We believe that TorchCP library will promote the adoption and advancement of conformal prediction within the deep learning community. In the future, TorchCP will be continuously updated with new algorithms, paradigms, and evaluation protocols. We also welcome any contributions for TorchCP.
Appendix A Details of conformal methods in TorchCP
@@ A.1 Split CP
Conformal prediction (Balasubramanian et al., 2014; Manokhin, 2022; Angelopoulos et al., 2023, 2024) is a statistical framework that generates uncertainty sets/intervals containing ground-truth label with a desired probability. Split Conformal Prediction is the most widely-used version of the conformal prediction procedure. Specifically, split conformal prediction divides a given dataset into two disjoint subsets: one for training the base classifier and the other for conformal calibration. We next outline the main process of split conformal prediction:
@@
    Train a deep learning model on the training dataset $\mathcal{D}_{tr}$, and define a non-conformity score function $V(\bm{x},y)$;
@@ A.2 Classification
Here, we consider multi-class classification with $K$ classes. Let $\mathcal{X}\subset\mathbb{R}^{d}$ be the input space and $\mathcal{Y}:=\{1,\dots,K\}$ be the label space. We use $(X,Y)\sim\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ to denote a random data pair satisfying a joint data distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ and $f:\mathcal{X}\rightarrow\mathbb{R}^{K}$ to denote a classification neural network. Thus, the classifier $\pi:\mathcal{X}\rightarrow\Delta^{K-1}$is defined as $\sigma\circ f$, where $\Delta^{K-1}$ is a ($K$-1)-dimensional probability simplex and $\sigma$ is a normalization function such as the softmax function. Let $\pi_{y}(x)$ be the probability of $y$ conditional on $x$, i.e., $P(y|x)$. In the following, we briefly describe methods of conformal prediction supported by TorchCP.
THR (Sadinle et al., 2019): The THR measures the similarity between the example and the data space by the conditional probability $P(Y|X)$: $V(\bm{x},y):=1-\pi_{y}(\bm{x})$.
@@
Class-wise CP (Vovk, 2012): Class-wise CP partitions the calibration data based on their labels $y$. The non-conformity scores’ quantiles are then computed separately for each class using the corresponding calibration subset.
Clustered CP (Ding et al., 2024): Clustered CP extends Class-wise CP by applying split conformal prediction to class clusters rather than individual classes. It achieves cluster-conditional coverage as an approximation of class-conditional guarantees.
@@
ConfTr (Stutz et al., 2022): ConfTr simulates the CP process based on end-to-end training by splitting mini-batches into calibration and prediction sets and designing differentiable conformal prediction steps: $\mathcal{L}_{\mathrm{ConfTr}}=\max\left(0,\sum_{k=1}^{K}\mathcal{C}_{\theta,k}(\bm{x})-\kappa\right),$ where $\theta$ is the trainable parameters of deep learning models and $\kappa$ is the target set size. The definition of $\mathcal{C}_{\theta,k}(\bm{x})$ is given by $\sigma((V(\bm{x},k)-\widehat{Q}_{1-\alpha})/\tau)$ where $\tau$ is a temperature hyperparameter and $\sigma$ is the sigmoid function.
ConfTS (Xi et al., 2024): ConfTs is a training method to optimize the $T$ of temperature scaling for improved efficiency of prediction sets. The objective of ConfTS is to minimize the efficiency gap, i.e., $\widehat{Q}_{1-\alpha}-V(\bm{x},y)$. Formally, the loss function of ConfTS is:

### v3

@@ Abstract
Conformal prediction (CP) is a robust statistical framework that generates prediction intervals or sets with guaranteed coverage probability, addressing the challenge of quantifying predictive uncertainty in deep learning. Despite advancements in deep learning architectures and datasets, reliable uncertainty estimation remains elusive, making CP increasingly vital. This paper introduces TorchCP, a PyTorch-native library designed to integrate state-of-the-art CP algorithms into deep learning tasks, including classification, regression, graph neural networks, and large language models. TorchCP offers a comprehensive suite of advanced methodologies, a modular design for easy customization, and full GPU-accelerated scalability. Released under the LGPL-3.0 license, TorchCP has gained widespread adoption with over 12,582 PyPi downloads. It is supported by approximately 16,132 lines of code, 564 unit tests achieving 100% coverage, and comprehensive documentation. By bridging statistics and computer science, TorchCP empowers researchers and practitioners to advance conformal prediction in diverse deep learning applications.
Keywords: Conformal Prediction, Deep Learning
@@ 1 Introduction
Conformal prediction (CP) is a powerful statistical framework generating prediction intervals or sets that contain true labels with a specified coverage probability (Vovk et al., 2005). In recent years, deep learning has evolved dramatically with advanced model architectures and vast datasets, yet it struggles to reliably measure predictive uncertainty, making conformal prediction an increasingly attractive solution. Beyond classification and regression, conformal prediction methods have been tailored for diverse deep learning tasks, such as graph neural networks (GNNs) for molecular modeling and large language models (LLMs) for natural language understanding. This underscores the need for a comprehensive and efficient CP toolbox tailored for deep learning models, facilitating robust and accessible conformal prediction methods across diverse deep learning applications.
This work presents TorchCP, a PyTorch-native library of conformal prediction for deep learning. Firstly, TorchCP implements a wide range of state-of-the-art CP algorithms encompassing both training and inference methods designed for deep learning tasks, e.g., classification (ordinal classification), regression (time series), GNN, and LLM. Secondly, the library’s modular design allows developers to easily use or customize each component (e.g., losses or scores) of the library. Lastly, TorchCP is fully built on PyTorch with GPU support and batch processing, enabling seamless scalability for large-scale deep learning tasks.
The TorchCP library, released under the LGPL-3.0 license, can be installed via PyPi with the command ‘‘pip install torchcp’’ and has been downloaded 12,582 times  [^1] https://pepy.tech/projects/torchcp. The implementation consists of approximately 16,132 lines of code and is thoroughly tested with 564 unit tests, achieving 100% test coverage. We also provide comprehensive online documentation  [^2] https://torchcp.readthedocs.io/en/latest/ to guide users through TorchCP’s features. We hope the library can bridge the gap between statistics and computer science, enabling researchers and practitioners from both backgrounds to advance conformal prediction in deep learning.
2 The library design and implementation
Figure 1: An illustration of the modular design of TorchCP.
2.1 Modular design
For low coupling, we separate the TorchCP library into separate modules that can be easily maintained, tested, and integrated. For each task, the code implementation of the library contains 3 modules: Trainer, Score, and Predictor. Specifically, Trainer defines training algorithms for improving conformal prediction, along with loss functions. Score defines non-conformity score functions that quantify how different a new sample is from the training data, e.g., LAC (Sadinle et al., 2019). Predictor defines CP workflows (e.g., calibration and prediction), such as Split CP  [^3] A detailed introduction of Split CP is presented in Appendix A.1 (Papadopoulos et al., 2002). The modular design allows a straightforward extension to new tasks or new CP methods that arise in this active research field.
We visualize the modular design of TorchCP in Figure 1, highlighting the relationships among the three modules. The Trainer module is used to train or fine-tune deep models, employed by the Predictor module for calibration and prediction. The Score module implements non-conformity score functions, enabling their flexible use across different trainers and predictors. This modular structure empowers users to independently select or customize trainers, score functions, and predictors for diverse conformal prediction tasks, significantly improving flexibility and maintainability. Additionally, TorchCP achieves strong extensibility by providing standardized module interfaces, making it straightforward to define new methods. Appendix B provides guidelines for adding customized modules using the provided interfaces, while Appendix E discusses the advantages of TorchCP ’s modular design.
2.2 Implementation
Table 1: The CP algorithms implemented in TorchCP.
Task Component Methods
Classification Trainer&Loss ConfTr (Stutz et al., 2022), ConfTS (Xi et al., 2025), C-Adapter (Liu et al., 2024), UA Loss (Einbinder et al., 2022), SCPO (Bellotti, 2021)
Score Function LAC (Sadinle et al., 2019), APS (Romano et al., 2020), RAPS (Angelopoulos et al., 2020), SAPS (Huang et al., 2024), KNN (Gammerman and Vovk, 2007), Margin (Löfström et al., 2015), TOPK (Angelopoulos et al., 2020)
Predictor Split CP (Papadopoulos et al., 2002) Class-conditional CP (Vovk, 2012), Clustered CP (Ding et al., 2024), Weighted CP (Tibshirani et al., 2019), RC3P (Shi et al., 2024)
Regression Score Function ABS (Papadopoulos et al., 2002), NORABS (Vovk et al., 2005), CQR (Romano et al., 2019), CQRR (Sesia and Candès, 2020), CQRM (Sesia and Candès, 2020), CQRFM (Kivaranovic et al., 2020), R2CCP (Guha et al., 2023)
Predictor Ensemble CP (Xu and Xie, 2021), ACI (Gibbs and Candes, 2021), AgACI (Zaffran et al., 2022)
GNN Trainer&Loss CF-GNN (Huang et al., 2023)
Score Function DAPS (Zargarbashi et al., 2023), SNAPS (Song et al., 2024)
Predictor NAPS (Clarkson, 2023)
LLM Predictor Conformal LLM (Quach et al., 2024)
Methodological scope
Table 1 outlines the CP algorithms included in TorchCP across 4 deep learning tasks: classification, regression (time series), GNN, and LLM , along with their corresponding evaluation metrics in Table 4. In total, TorchCP provides 6 training algorithms, 16 score functions, and 10 predictors tailored to various tasks. A brief overview of all supported methods and their advantages is available in Appendix A.
Batch Processing and GPU Support
TorchCP is a PyTorch-native library equipped with GPU acceleration and batch processing capabilities. It enables efficient computation on large datasets and scales seamlessly to real-world deep learning applications. Appendix D.2 presents empirical evidence of its efficiency advantages on large-scale datasets.
Tutorials
TorchCP aims to provide a user-friendly framework for conformal prediction in deep learning. The library contains illustrative code showcasing functionality for common tasks (See the “examples” folder  [^4] https://github.com/ml-stat-Sustech/TorchCP/tree/master/examples), and Appendix C offers a practical guide to help users identify relevant examples. In Appendix D, we provide reproduced results of classical CP methods, confirming the correctness of our implementation. Below, we present a generic example  [^5] The full code of the example is available at https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/classification_conftr_cifar100.py to show the usability of TorchCP. First, we employ ConfTr (Stutz et al., 2022) to train a model for conformal prediction (optional, using standard trained models also work):
1from torchcp.classification.trainer import ConfTrTrainer
2trainer = ConfTrTrainer(init_model, alpha=0.1) # Create a trainer with a NN
3trained_model = trainer.train(train_loader, num_epochs=10) # Training 10 epochs
With the trained model, users can proceed with conformal prediction by selecting an appropriate score function and predictor. As an example, we implement Split CP with a miscoverage rate of $\alpha=0.1$ using the LAC score :
1predictor = SplitPredictor(score_function=LAC(), model=trained_model, alpha=0.1)
2predictor.calibrate(cal_loader) # Compute the threshold on calibrate data
3predict_set = predictor.predict(test_instance) # Predict the set for an instance
3 Related library
Conformal Prediction (CP) is a distribution-free framework for quantifying the uncertainty of machine learning model predictions. While some uncertainty quantification libraries, such as PUNCC (Mendil et al., 2023), implement certain CP algorithms, their limited efficiency, scalability, and algorithmic coverage restrict their use in deep learning. For instance, most libraries support only a narrow range of tasks, such as regression and classification, and lack support for complex applications like generative modeling or time-series forecasting. Additionally, the absence of GPU acceleration and batch processing in these libraries makes them inefficient for large-scale scenarios. In contrast, TorchCP offers a clean interface, enhanced extensibility, high efficiency, and a comprehensive set of CP algorithms. For example, TorchCP reduces inference time by up to 90% compared to PUNCC on large-scale datasets (see Appendix D.2 for details). We summarize the comparison of different libraries in Table 2 and discuss the advantages of TorchCP ’s modular design in Appendix E.
Table 2: Feature comparison of open-source libraries supporting conformal prediction. “Training” refers to CP-specific training algorithms, and “Online” denotes online CP.
Library Task Support Functionality Support Backend
Classification Regression GNN LLM Training Online GPU Batch
nonconformist (Linusson et al., 2020) Numpy
TorchUQ (Zhao et al., 2024) PyTorch
@@ 4 Conclusion and Future work
In this work, we introduced TorchCP, a PyTorch-native library that seamlessly integrates conformal prediction (CP) into deep learning workflows. Supporting a wide range of state-of-the-art CP algorithms, TorchCP leverages PyTorch’s GPU acceleration and batch processing, offering a modular design for easy customization and user-friendly documentation for accessibility. Looking ahead, we plan to expand TorchCP by incorporating emerging CP algorithms for evolving architectures like diffusion models and multimodal models, while optimizing for real-time performance. Through open-source collaboration on GitHub, we aim to foster community-driven innovation, empowering reliable AI in fields such as healthcare, finance, and autonomous systems.
Appendix A Details of Conformal Prediction Methods in TorchCP
In this section, we provide brief introductions to the methods implemented in TorchCP. Furthermore, Table 3 offers practical guidelines to help users select appropriate methods based on their specific application needs.
Table 3: Advantages or use cases of CP methods across different tasks.
Component Method Advantage/Use case
Classification
Score LAC To achieve the smallest prediction sets
APS To approximate conditional coverage
RAPS For smaller, more stable sets than APS
TOPK A simple baseline with only logit ranking
SAPS Improving efficiency, conveying instance-wise uncertainty
Margin A non-conformity function based on prediction margin
KNN A score function based solely on feature-space nearest neighbors without a predictive model.
Predictor Class-conditional CP To achieve coverage for every class
Clustered CP For class-conditional coverage with limited data per class
Weighted CP Conformal prediction under covariate shift
Loss & Trainer ConfTr Efficient prediction sets with conformal-aware loss
C-Adapter Optimizing efficiency based on $\alpha$-free loss
ConfTS Improving efficiency via temperature scaling
Regression
Score ABS Simple and interpretable residual-based scores
NORABS Improving robustness of prediction intervals via noise-aware normalization
CQR Flexible interval width via quantile regression
R2CCP Shorter prediction intervals than CQR via recasting regression as classification
Predictor ACI Reliable coverage under distributional shifts
AgACI Aggregates ACI predictions with adaptive weighting
Ensemble CP Stable prediction intervals without data splitting via model ensembling
GNN
Score DAPS Improves efficiency of APS using structural neighbors
SNAPS Extends DAPS with feature neighbors for better efficiency
NAPS Constructing prediction sets for inductive node classification via Weighted CP
Loss & Trainer CF-GNN Reducing set size by attaching an adapter with a topology-aware inefficiency loss
LLM
Predictor Conformal LLM To achieve statistically guaranteed prediction sets for language models through calibrated sampling
@@ A.1 Split CP
Conformal prediction (Papadopoulos et al., 2002; Balasubramanian et al., 2014; Manokhin, 2022; Angelopoulos et al., 2023, 2024) is a statistical framework that generates uncertainty sets/intervals containing ground-truth labels with a desired probability. Split Conformal Prediction (Papadopoulos et al., 2002) (Split CP) is the most widely-used version of the conformal prediction procedure. Specifically, Split CP divides a given dataset into two disjoint subsets: one for training the base model and the other for conformal calibration. We next outline the main process of split conformal prediction:
@@
    Train a model on the training dataset $\mathcal{D}_{tr}$, and then define a non-conformity score function $V(\bm{x},y)$;
@@ A.2 Classification
Here, we provide a brief overview of conformal prediction methods for multi-class classification tasks as supported by TorchCP. Let $\mathcal{X}\subset\mathbb{R}^{d}$ be the input space and $\mathcal{Y}:=\{1,\dots,K\}$ be the label space. We use $(X,Y)\sim\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ to denote a random data pair satisfying a joint data distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ and $f:\mathcal{X}\rightarrow\mathbb{R}^{K}$ to denote a classification neural network. Thus, the classifier $\pi:\mathcal{X}\rightarrow\Delta^{K-1}$is defined as $\sigma\circ f$, where $\Delta^{K-1}$ is a ($K$-1)-dimensional probability simplex and $\sigma$ is a normalization function such as the softmax function. Let $\pi_{y}(x)$ be the probability of $y$ conditional on $x$, i.e., $P(y|x)$.
LAC (Sadinle et al., 2019): The LAC measures the similarity between the example and the data space by the conditional probability $P(Y|X)$: $V(\bm{x},y):=1-\pi_{y}(\bm{x})$.
@@
Class-conditional CP (Vovk, 2012): Class-conditional CP partitions the calibration data based on their labels $y$. The non-conformity scores’ quantiles are then computed separately for each class using the corresponding calibration subset.
Clustered CP (Ding et al., 2024): Clustered CP extends Class-conditional CP by applying split conformal prediction to class clusters rather than individual classes. It achieves cluster-conditional coverage as an approximation of class-conditional guarantees.
@@
RC3P (Shi et al., 2024): RC3P addresses the inefficiency of class-conditional conformal prediction, which often results in large prediction sets, especially for imbalanced classification tasks. The core idea is to selectively perform class-wise thresholding only for labels with low top-k error, by integrating a label rank calibration step into standard class-wise conformal calibration.
ConfTr (Stutz et al., 2022): ConfTr simulates the CP process based on end-to-end training by splitting mini-batches into calibration and prediction sets and designing differentiable conformal prediction steps: $\mathcal{L}_{\mathrm{ConfTr}}=\max\left(0,\sum_{k=1}^{K}\mathcal{C}_{\theta,k}(\bm{x})-\kappa\right),$ where $\theta$ is the trainable parameters of deep learning models and $\kappa$ is the target set size. The definition of $\mathcal{C}_{\theta,y}(\bm{x})$ is given by $\sigma((\widehat{Q}_{1-\alpha}-V(\bm{x},y))/\tau)$ where $\tau$ is a temperature hyperparameter and $\sigma$ is the sigmoid function.
ConfTS (Xi et al., 2025): ConfTs is a training method to optimize the $T$ of temperature scaling for improved efficiency of prediction sets. The objective of ConfTS is to minimize the efficiency gap, i.e., $\widehat{Q}_{1-\alpha}-V(\bm{x},y)$. Formally, the loss function of ConfTS is:
@@
SCPO (Bellotti, 2021): SCPO proposes an approach to train the deep learning model directly by optimizing for maximum predictive efficiency. To enable end-to-end training, the conformal predictor is approximated by a differentiable surrogate objective. Gradient descent is then used to minimize the following loss:
$\mathcal{L}_{\mathrm{SCPO}}=\sum_{k=1}^{K}\mathcal{C}_{\theta,k}(\bm{x})+\lambda\left(\mathcal{C}_{\theta,y}(\bm{x})-(1-\alpha)\right)^{2},$
where $\lambda>0$ controls the relative importance of the two objective components.
Uncertainty-aware (UA) Loss (Einbinder et al., 2022): Uncertainty-aware Loss introduces a novel training objective that integrates conformal calibration directly into model training by encouraging the conformity scores to follow a uniform distribution. The novel uncertainty-aware component of this loss is:
$\mathcal{L}_{u}=\sup_{w\in[0,1]}\left|\mathbb{E}[\mathds{1}(V(\bm{x},y,u)-w]\right|,$
where $V(\bm{x},y,u):=\sum\limits_{y^{\prime}\in\mathcal{Y}}\pi_{y^{\prime}}(\bm{x})\mathds{1}(r_{f}(\bm{x},y^{\prime})<r_{f}(\bm{x},y))+u\cdot\pi_{y}(\bm{x})$.